Electromagnetic radiation having a 14.0 mu wavelength is classified as infrared radiation. What is it's frequency?
1
Expert's answer
2016-02-12T00:01:01-0500
We will use formula: c=λ*ν,where λ is wavelength (given 12 mu=12*〖10〗^(-6) meters)and ν is frequency.
c is speed of electromagnetic wave (c=3*〖10〗^8 m/sec)
So, ν=c/λ=(3*〖10〗^8)/(12*〖10〗^(-6) )=0.25*〖10〗^14 Hz=25 THz

Answer v=25 THz